a satellite is powered by a small nuclear reactor that puts out 15w.how much mass is converted into energy in 10years life span of generator

Grade:12

1 Answers

SWATI RANE
5 years ago
Energy=15(j/s)×10(y)×365.25(d/y)×24(h/d)×360
0(s/h)
=4.73×10^9 J
But E = mc^2
mc^2=4.73×10^9
m=4.73×10^9/(3×10^8)^2
m=5.26×10^-8 kg
m=52.6 micro g